# Dynamics of An Information Theoretic Analog of Two Masses on a Spring

### v1

@@ Abstract
In this letter we investigate an information theoretic analogue of the classic two masses on spring system, arising from a physical interpretation of Friston’s free energy principle in the theory of learning in a system of agents. Using methods from classical mechanics on manifolds, we define a kinetic energy term using the Fisher metric on distributions and a potential energy function defined in terms of stress on the agents’ beliefs. The resulting Lagrangian (Hamiltonian) produces a variation of the classic DeGroot dynamics. In the two agent case, the potential function is defined using the Jeffrey’s divergence and the resulting dynamics are characterized by a non-linear spring. These dynamics produce trajectories that resemble flows on tori but are shown numerically to produce chaos near the boundary of the space. We then investigate persuasion as an information theoretic control problem where analysis indicates that manipulating peer pressure with a fixed target is a more stable approach to altering an agent’s belief than providing a slowly changing belief state that approaches the target.
preprint: APS/123-QED
@@
In this letter, we rephrase the free energy principle in terms of free entropy, which arises naturally from the information geometry of the Fisher metric [16, 17, 18]. This quantity and the corresponding techniques from information geometry have been used extensively in statistical physics [19, 20, 21, 22, 23, 24, 25, 26, 27].
@@
Let $p(\mathbf{x}|\bm{\eta})$ be a probability distribution parameterized by $\bm{\eta}$. For two distributions $p(\mathbf{x}|\bm{\eta})$ and $q(\mathbf{x}|\bm{\eta})$, the Kullback-Liebler divergence is given by,
@@
Unlike Euclidean spaces, statistical manifolds admit divergences, which are locally consistent with the Fisher metric, but are themselves not true distances. The Kullbach-Liebler divergence, already introduced, is one such divergence. In particular, we can locally write,
$D_{KL}\left[p(\mathbf{x}|\bm{\eta})|p(\mathbf{x}|\bm{\eta}_{0})\right]=\\ \frac{1}{2}\mathbf{g}_{jk}(\bm{\eta}_{0})\Delta\eta_{0}^{j}\Delta\eta_{0}^{k}+O(\left\lVert\bm{\Delta\eta}_{0}\right\rVert^{3}),$
showing the relationship between the metric and the divergence [57].
@@
For the remainder of this paper, we consider the setting where $q$ is the parameter in a Bernoulli random variable representing an agent’s belief about a certain outcome or opinion about a certain bivalent topic. If $q^{\prime}\in(0,1)$ is a fixed belief point (i.e., an external reference point) then the Kullbach-Liebler divergence yields an information theoretic “squared distance” from $q$ to $q^{\prime}$ as,
@@
$\tilde{L}(q,q^{\prime})=e^{-\alpha t}\left[\frac{m}{2}\frac{\dot{q}^{2}}{q(1-q)}-\frac{k}{2}D(q,q^{\prime})^{2}\right].$
@@
$\tilde{H}(q,q^{\prime})=e^{-\alpha t}\frac{q(1-q)}{2m}p_{i}^{2}+e^{\alpha t}\frac{k}{2}D(q,q^{\prime})^{2}.$
@@
where $H_{0}$ is an integrable Hamiltonian, $\epsilon$ is a perturbation constant and $H_{1}$ is the nonlinear Hamiltonian perturbation. Since the coordinate transform in Eq. 10 is a diffeomorphism, we may expect to see Hamiltonian chaos arise in the dynamics generated by Eq. 9. We find evidence for weak chaos [62] associated to behavior near the manifold boundary. This is shown in Fig. 3, where we see a decaying estimate for the maximum Lyapunov exponent, followed by a slight increase, using the Wolf algorithm [63]. In Fig. 3, $k=2$, which improves numerical stability of the long trajectory required to execute the Wolf algorithm. The fact that the maximum Lyapunov exponent is positive and the dynamics play out on a compact set is sufficient to conclude that they are (weakly) chaotic.
@@
Figure 3: The Wolf algorithm [63] is used to estimate the Lyapunov spectrum for the Hamiltonian dynamics arising from Eq. 9. The results suggest weak chaos caused by behavior near the boundaries of the space.
@@
This is consistent with prior results on Hamiltonian chaos on compact sets defined by the unit simplex that showed that the chaotic dynamics arose as a result of interactions with the boundary [64]. This Hamiltonian system is unusual in that it has an infinite line of fixed points $q_{1}=q_{2}$ and $p_{1}=p_{2}=0$. The result is a degeneracy, which makes application of the classical KAM theorem impossible [65] because the dynamics generated by $H_{0}$ evolve on a cylinder homeomorphic to $\mathbb{S}^{1}\times\mathbb{R}$, rather than the two-torus $\mathbb{S}^{2}$. From this we see that sensitivity to initial conditions and the classical behaviors associated to weak chaos emerge as a result of asymmetries in the initial conditions. This is shown in Fig. 4, where we fix $q_{2}(0)=\tfrac{1}{2}$ and vary $q_{1}(0)$, illustrating the distinct trajectory structures that emerge.
Figure 4: As $q_{1}(0)$ is varied, the structure of the trajectory $(q_{1},q_{2})$ changes from an ellipse to a complex curve.
@@
Despite the results of the Wolf algorithm, the trajectories in Fig. 4 seem well-behaved. This is because of the region in the interior of unit simplex, where the geometry is locally Euclidean. In this region, the behavior of the system is given almost entirely by a two-mass and spring system. Non-integrability (and thus the destruction of additional conserved quantities) is only introduced when the dynamics interact with the boundary, which induces an effective nonlinear force, not present in Euclidean space.
@@
$\mathcal{O}=\min\int_{t_{0}}^{t_{f}}\frac{p^{2}}{2}+Cdt,$ (13)
where $C$ is a cost associated to the control parameter. We include the kinetic energy term in the performance index because without some cost associated to momentum, Eq. 13 degenerates into a “bang-bang” style control problem [67]. Additionally, the momentum cost is related to the “accumulated surprise” generated by the control signal [66], and minimizing the accumulated surprise reduces the chances that persuasive actions will be completely ignored. We work in $\theta$ coordinates for simplicity of form and because the resulting dynamics are more numerically stable.
@@
$C_{D}=\beta D[\cos^{2}(\theta/2),\cos^{2}(\theta^{\prime}/2)],$
and suppose that the state dynamics are given by the full Hamiltonian $H$ from Eq. 11 with $\theta_{1}=\theta$ and $\theta_{2}=\theta^{\prime}$. The control Hamiltonian (which is distinct from the Hamiltonian producing the system dynamics) [68] is,
$\mathcal{H}=\frac{p^{2}}{2}+C_{D}+\lambda\frac{\partial H}{\partial p}-\mu\frac{\partial H}{\partial\theta},$ (14)
where $\lambda$ and $\mu$ are the Lagrange multipliers associated to $\theta$ and $p$, respectively. The optimal control path is characterized by Pontryagin’s maximum principle [69], as usual. Since the resulting system of differential equations is poorly behaved, we also consider the second-order approximation to $\mathcal{H}$ with cost function,
$C_{S}=\frac{\beta}{4}(\theta-\theta^{\prime})^{2},$
@@
As long as both belief boundary conditions are sufficiently removed from the edges of the space, the $\theta^{\prime}$-control problem is well behaved and the second-order approximation is reasonably accurate. The system becomes poorly behaved numerically in regions near certainty because the curvature of the information manifold makes both the cost and the state dynamics stiff. Additionally, the approximate control problem begins to exhibit wrapping in this regime, resulting in non-physical solutions. Since the control variable acts via the KL-divergence, which is unbounded on the manifold, it would seem it is possible to reach any target belief within an arbitrary time horizon. However the Fisher-Rao metric is poorly conditioned near the manifold boundaries so that in practice the control problem becomes intractable, in addition to a potential breakdown in overall model fidelity.
To demonstrate, consider boundary values $\theta(0)=2\arccos\sqrt{0.5}$ and $\theta(3)=2\arccos\sqrt{0.9}$ ($q_{0}=0.5$ and $q_{f}=0.9$, respectively) and parameters $\alpha=0.5$, $\beta=10$, and $k=1$. Additionally assume zero initial momentum, $p(0)=0,$ and the transversality boundary value $\mu(3)=0$. For the exact formulation of the control problem, a globally optimal solution can be found numerically using standard solvers. For the approximate formulation, an approximately optimal control path can be written analytically, because the problem is a LQC problem [68]. We can model the behavior of this solution under the full system dynamics by using the approximate control to propagate the forward initial value problem defined via $H$. Fig. 5 shows the resulting the control paths $\theta^{\prime}$ and corresponding belief evolution in probability coordinates $q$ for both the exact and approximate control problems. The two solutions are fairly similar with the approximate solution achieving a final belief value of $q_{f}=0.93$ instead of the intended $q_{f}=0.9$.
@@
While proximity to the geometric boundary drives stiffness in the control problem, this can be ameliorated to some extent by choosing a longer time horizon $T$. We can very roughly approximate a lower bound for an acceptable time horizon by solving the approximate control problem with the maximum (minimum) possible control input $\theta^{\prime}=\pi$ ($\theta^{\prime}=0$), to estimate the shortest horizon over which the desired belief target can be met without “wrapping” the control. While this lower bound doesn’t have a clean analytic formula, it is straightforward to compute numerically. Using the parameter values above the minimum time horizon to avoid wrapping is $T\approx 1.9$. Numerical experimentation indicates that reasonable control solutions can be found for $T>2.2$, below which both the exact system and the forward simulation of the approximate control system become stiff. Notably as $T$ approaches this threshold the approximate control curve gets closer to the boundaries of the geometry, and the approximation breaks down.
As an alternative to controlling the belief distribution by changing $\theta^{\prime}$, we can instead fix $\theta^{\prime}=\theta_{f}$ to be the desired belief and use the spring constant $k$ as the control parameter. We use the same objective function as Eq. 13 with the quadratic cost $C=\beta k^{2}$. We use this cost for both the exact dynamics, given by $H$, and approximate dynamics, given by $H_{0}$. The resulting $k$-control problems do not appear to have a simple closed-form solution. However, the approximate control problem is again a LQC problem and both formulations are much better behaved numerically than the $\theta^{\prime}$-control formulation. This is because the control parameter $k$ does not lie on the information manifold and is not as strongly impacted by the manifold curvature.
@@
While a more thorough analysis of this information control problem, including joint $(k,\theta^{\prime})$-control, is outside the scope of this letter, this initial analysis indicates that the spring constant (i.e., social pressure) is a more robust control parameter than the observed belief. At a high level this is because the spring constant produces information potentials which scale linearly, while the observed belief can only achieve large potentials by interacting with the highly curved portions of the manifold boundary. Additionally, the simpler second-order approximation to the control problem, which utilizes traditional spring potentials and is a LQC problem, can produce effective control solutions, particularly in regions of high entropy.
In this letter, we introduced a model of belief combining Friston’s minimum energy principle and information geometry, making use of the Bayesian brain hypothesis. Using a Bernoulli distribution as a model of an agent’s belief about a bivalent topic (e.g., humans cause climate change), we showed how to recover a variant of DeGroot opinion dynamics via a classical Lagrangian. We further showed conditions under which opinion converges or comes to consensus through the use of a dissipative Lagrangian. In studying the nonlinear dynamics for two agents arising from the non-dissipative form of the Lagrangian, we showed that the corresponding Hamiltonian was a perturbation of a well-known integrable Hamiltonian. Despite the fact that the KAM theorem does not apply in this case, we illustrated symplectic flow near the fixed points and used Wolf’s algorithm to numerically prove the existence of chaotic behavior in trajectories near the boundary.

### v2

@@ Abstract
In this short communication we investigate an information theoretic analogue of the classic two masses on spring system, arising from a physical interpretation of Friston’s free energy principle in the theory of learning in a system of agents. Using methods from classical mechanics on manifolds, we define a kinetic energy term using the Fisher metric on distributions and a potential energy function defined in terms of stress on the agents’ beliefs. The resulting Lagrangian (Hamiltonian) produces a variation of the classic DeGroot dynamics. In the two agent case, the potential function is defined using the Jeffrey’s divergence and the resulting dynamics are characterized by a non-linear spring. These dynamics produce trajectories that resemble flows on tori but are shown numerically to produce chaos near the boundary of the space. We then investigate persuasion as an information theoretic control problem where analysis indicates that manipulating peer pressure with a fixed target is a more stable approach to altering an agent’s belief than providing a slowly changing belief state that approaches the target.
I Introduction and Background
@@
In this short communication, we rephrase the free energy principle in terms of free entropy, which arises naturally from the information geometry of the Fisher metric [16, 17, 18]. This quantity and the corresponding techniques from information geometry have been used extensively in statistical physics [19, 20, 21, 22, 23, 24, 25, 26, 27].
@@
Let $p(\mathbf{x}|\bm{\eta})$ be a probability distribution parameterized by $\bm{\eta}$. For two distributions $p(\mathbf{x}|\bm{\eta})$ and $q(\mathbf{x}|\bm{\eta})$, the Kullback-Liebler (KL) divergence is given by,
@@
In addition to the usual Riemannian geodesic distance, statistical manifolds admit divergences, which are locally consistent with the Fisher metric and yield affine connections [58, 57], but are themselves not true distances. The KL-divergence, already introduced, is one such divergence. In particular, we can locally write,
$D_{KL}\left(p(\mathbf{x}|\bm{\eta})|p(\mathbf{x}|\bm{\eta}_{0})\right)=\sum_{j,k}\frac{1}{2}\mathbf{g}_{jk}(\bm{\eta}_{0})\Delta\eta^{j}\Delta\eta^{k}+O(\left\lVert\bm{\Delta\eta}\right\rVert^{3}),$
showing the relationship between the metric and the divergence [57]. In this communication we use the Fisher metric to measure distance and length locally, while the KL-divergence will play the role of a global Mahalanobis distance.
@@ II Opinion Dynamics as a Nonlinear Spring System
For the remainder of this paper, we consider the setting where $q$ is the parameter in a Bernoulli random variable representing an agent’s belief about a certain outcome or opinion about a certain bivalent topic. If $q^{\prime}\in(0,1)$ is a fixed belief point (i.e., an external reference point) then the KL-divergence yields an information theoretic “squared distance” from $q$ to $q^{\prime}$ as,
@@
$\tilde{L}(q,q^{\prime})=e^{-\alpha t}\left[\frac{m}{2}\frac{\dot{q}^{2}}{q(1-q)}-\frac{k}{2}D(q,q^{\prime})\right].$
@@
$\tilde{H}(q,q^{\prime})=e^{-\alpha t}\frac{q(1-q)}{2m}p_{i}^{2}+e^{\alpha t}\frac{k}{2}D(q,q^{\prime}).$
@@ III Chaotic Dynamics in the Two Agent System
where $H_{0}$ is an integrable Hamiltonian, $\epsilon$ is a perturbation constant and $H_{1}$ is the nonlinear Hamiltonian perturbation. Since the coordinate transform in Eq. 10 is a diffeomorphism, we expect to see Hamiltonian chaos arise in the dynamics generated by Eq. 9. We find evidence for weak chaos [63] associated to behavior near the manifold boundary. To obtain this result, we used a time-varying estimate for the maximum Lyapunov exponent of the dynamical system using the Wolf algorithm [64]. In particular, we used Sandri’s implementation of the Wolf algorithm [65], which returns an estimation of the the Lyapunov spectrum using a single long trajectory after an initial burn-in period. Fig. 3 shows the instantaneous estimates of the maximum Lyapunov exponent during algorithm execution over the long trajectory. The final estimate provided by the algorithm for the maximum Lyapunov exponent is $\lambda\approx 0.0003>0$. Positivity of the maximum Lyapunov exponent is further supported by the fact that the trajectory shown in Fig. 3 turns up after an initial period of decay. In Fig. 3, the spring constant $k=2$, which improves numerical stability of the long trajectory required to execute the Wolf algorithm. The fact that the maximum Lyapunov exponent is positive and the dynamics play out on a compact set is sufficient to conclude that they are (weakly) chaotic.
@@
Figure 3: The Wolf algorithm [64] is used to estimate the Lyapunov spectrum for the Hamiltonian dynamics arising from Eq. 9. The upturn in the plot of the estimate for the maximum Lyapunov exponent suggests weak chaos caused by behavior near the boundaries of the space. The Wolf algorithm estimate for the maximum Lyapunov is $\lambda\approx 0.0003>0$.
@@
This is consistent with prior results for Hamiltonian chaos on a compact set defined by the unit simplex that showed that the chaotic dynamics arose as a result of interactions with the boundary [66]. This Hamiltonian system is unusual in that it has an infinite line of fixed points $q_{1}=q_{2}$ and $p_{1}=p_{2}=0$. The result is a degeneracy, which makes application of the classical Kolmogorov–Arnold–Moser (KAM) theorem impossible [67] because the dynamics generated by $H_{0}$ evolve on a cylinder homeomorphic to $\mathbb{S}^{1}\times\mathbb{R}$, rather than the two-torus $\mathbb{S}^{2}$. From this we see that sensitivity to initial conditions and the classical behaviors associated to weak chaos emerge as a result of asymmetries in the initial conditions. This is shown in Fig. 4, where we fix $q_{2}(0)=\tfrac{1}{4}$ and vary $q_{1}(0)$, illustrating the distinct trajectory structures that emerge. In Fig. 4, we project all trajectories onto the plane $q_{1}+q_{2}=1$, effectively collapsing the cylinders on which the dynamics evolve near the interior fixed point $q_{1}=q_{2}=\tfrac{1}{2}$ into ellipses.
Figure 4: Projection of solutions to Eq. 11 to the plane spanned by $u=\left\langle{0,0,1}\right\rangle$ and $v=\langle\frac{1}{\sqrt{2}},-\frac{1}{\sqrt{2}},0\rangle$. We fix $q_{2}(0)=\tfrac{1}{4}$ and vary $q_{1}(0)$. As the initial condition for $q_{1}$ is varied, the structure of the trajectory $(q_{1},q_{2})$ changes. When $q_{1}(0)$ is close to $\tfrac{1}{2}$, the dynamics are in a spatial region that is largely flat and the trajectories are projected ellipses. When $q_{1}(0)$ is far from $\tfrac{1}{2}$, the geometry influences the trajectory and the trajectories are more chaotic.
@@
These nested cylinders are (essentially) degenerate KAM tori. This is because of the region in the interior of unit simplex, where the geometry is locally Euclidean. In this region, the behavior of the system is given almost entirely by a two-mass and spring system. Non-integrability (and thus the destruction of additional conserved quantities) is only introduced when the dynamics interact with the boundary, which induces an effective nonlinear force, not present in Euclidean space. That is, when $q_{1}(0)$ is near the boundary, we see a more chaotic trajectory emerge in Fig. 4 a result of the boundary interactions.
IV Control of Opinion Dynamics
@@
$\mathcal{O}=\int_{t_{0}}^{t_{f}}\frac{p^{2}}{2}+C(k,\theta,\theta^{\prime})\,dt,$ (13)
where $C(k,\theta,\theta^{\prime})$ is a cost associated to the control parameter. The exact definition of $C$ depends on the method of control (external belief or spring constant) and is specified in the following. We include the kinetic energy term in Eq. 13 because without some cost associated to momentum, Eq. 13 degenerates into a “bang-bang” style control problem [69]. Additionally, the momentum cost is related to the “accumulated surprise” generated by the control signal [68], and minimizing the accumulated surprise is consistent with the free energy principle.
We work in $\theta$ coordinates for simplicity of form and because the resulting dynamics are more numerically stable. The associated control Hamiltonian (which is distinct from the system Hamiltonian $H$) [70] is,
$\mathcal{H}=\frac{p^{2}}{2}+C(k,\theta,\theta^{\prime})+\lambda\frac{\partial H}{\partial p}-\mu\frac{\partial H}{\partial\theta},$ (14)
where $\lambda$ and $\mu$ are the Lagrange multipliers associated to $\theta$ and $p$, respectively.
@@
$C_{D}(\theta,\theta^{\prime})=\beta D[\cos^{2}(\theta/2),\cos^{2}(\theta^{\prime}/2)],$
and suppose that the state dynamics are derived from the full Hamiltonian $H$ from Eq. 11 with $\theta_{1}=\theta$ and $\theta_{2}=\theta^{\prime}$. The optimal control path resulting from Eq. 14 is characterized by Pontryagin’s maximum principle [71], as usual. Since the resulting system of differential equations is poorly behaved, we also consider the second-order approximation to $\mathcal{H}$ with cost function,
$C_{S}(\theta,\theta^{\prime})=\frac{\beta}{4}(\theta-\theta^{\prime})^{2},$
@@
As long as both belief boundary conditions are sufficiently removed from the edges of the space, the $\theta^{\prime}$-control problem is well behaved and the second-order approximation is reasonably accurate. The system becomes poorly behaved numerically in regions near certainty because the curvature of the information manifold makes both the cost and the state dynamics stiff. Additionally, the approximate control problem begins to exhibit wrapping in this regime, resulting in non-physical solutions. Since the control variable acts via the KL-divergence, which is unbounded on the manifold, it would seem possible to reach any target belief within an arbitrary time horizon. However the Fisher-Rao metric is poorly conditioned near the manifold boundaries so that in practice the control problem becomes intractable, in addition to a potential breakdown in overall model fidelity.
To demonstrate, consider boundary values $\theta(0)=2\arccos\sqrt{0.5}$ and $\theta(3)=2\arccos\sqrt{0.9}$ ($q_{0}=0.5$ and $q_{f}=0.9$, respectively) and parameters $\alpha=0.5$, $\beta=10$, and $k=1$. Additionally assume zero initial momentum, $p(0)=0,$ and the transversality boundary value $\mu(3)=0$. For the exact formulation of the control problem, a globally optimal solution can be found numerically using standard solvers. For the approximate formulation, an approximately optimal control path can be written analytically, as it is a LQC problem [70]. We can model the behavior of this solution under the full system dynamics by using the approximate control to propagate the forward initial value problem defined via $H$. Fig. 5 shows the resulting control paths $\theta^{\prime}$ and corresponding belief evolution in probability coordinates $q$ for both the exact and approximate control problems. The two solutions are fairly similar with the approximate solution achieving a final belief value of $q_{f}=0.93$ instead of the intended $q_{f}=0.9$.
@@
While proximity to the geometric boundary drives stiffness in the control problem, this can be ameliorated to some extent by choosing a longer time horizon $T$. We can very roughly approximate a lower bound for an acceptable time horizon by solving the approximate control problem with the maximum (minimum) possible control input $\theta^{\prime}=\pi$ ($\theta^{\prime}=0$), to estimate the shortest horizon over which the desired belief target can be met without “wrapping” the control. While this lower bound doesn’t have a clean analytic formula, it is straightforward to compute numerically. Using the parameter values above the minimum time horizon to avoid wrapping is $T\approx 1.9$. Numerical experimentation indicates that reasonable control solutions can be found for $T>2.2$, below which both the exact system and the forward simulation of the approximate control system become stiff. As $T$ approaches this threshold the approximate control curve gets closer to the boundaries of the geometry, and the approximation breaks down.
As an alternative to controlling the belief distribution by changing $\theta^{\prime}$, we can instead fix $\theta^{\prime}=\theta_{f}$ to be the desired belief and use the spring constant $k$ as the control parameter. We use the same objective function as Eq. 13 with the quadratic cost $C_{K}(k)=\beta k^{2}$. We use this cost for both the exact dynamics, given by $H$, and approximate dynamics, given by $H_{0}$. The resulting $k$-control problems do not appear to have a simple closed-form solution. However, the approximate control problem is again a LQC problem and both formulations are much better behaved numerically than the $\theta^{\prime}$-control formulation. This is because the control parameter $k$ does not lie on the information manifold and is not as strongly impacted by the manifold curvature.
@@
While a more thorough analysis of this information control problem, including joint $(k,\theta^{\prime})$-control, is outside the scope of this paper, this initial analysis indicates that the spring constant (i.e., social pressure) is a more robust control parameter than the observed belief. At a high level this is because the spring constant produces information potentials which scale linearly, while the observed belief can only achieve large potentials by interacting with the highly curved portions of the manifold boundary. Additionally, the simpler second-order approximation to the control problem, which utilizes traditional spring potentials and is a LQC problem, can produce effective control solutions, particularly in regions of high entropy.
V Conclusion and Future Directions
In this paper, we introduced a model of belief combining Friston’s minimum energy principle and information geometry, making use of the Bayesian brain hypothesis. Using a Bernoulli distribution as a model of an agent’s belief about a bivalent topic (e.g., humans cause climate change), we showed how to recover a variant of DeGroot opinion dynamics via a classical Lagrangian. We further showed conditions under which opinion converges or comes to consensus through the use of a dissipative Lagrangian. In studying the nonlinear dynamics for two agents arising from the non-dissipative form of the Lagrangian, we showed that the corresponding Hamiltonian was a perturbation of a well-known integrable Hamiltonian. Despite the fact that the KAM theorem does not apply in this case, we illustrated symplectic flow near the fixed points and used Wolf’s algorithm to numerically prove the existence of chaotic behavior in trajectories near the boundary.